# Either a Confidence Interval Covers, or It Doesn’t (Or Does It?): A Model-Based View of Ex-Post Coverage Probability

### 05 Mar 2026 v3

@@
Disclaimer: The findings and conclusions in this report are those of the author and do not necessarily represent the official position of the Centers for Disease Control and Prevention.
@@ 3.1 Confidence intervals and coverage indicators
3. 2 Infinite sequences of experiments
@@
3.3 A small aside with Borel-Cantelli
As a small aside, it may be instructive to recast the infinite sequence $(I_{1},I_{2},\ldots)$ above as sampling with replacement from the “population” of intervals that our chosen procedure can produce from our data. When we are working in a discrete or finite–population setting–for example, if our data take values on a finite grid because of limited measurement resolution (the case for most, if not all physical experiments), or we are literally sampling from a finite list of outcomes–there will be only countably many distinct interval values $J$ that can appear, and some of them will occur with strictly positive probability
$\pi_{\theta}(J)=P_{\theta}\{I(X)=J\}>0.$
Now imagine picking one such interval value $J$ (with fixed numerical endpoints). It is true that every time we run our procedure, we will either see $J$ or not. Importantly, though, because the events $\{I_{i}=J\}$ are independent across all runs of the procedure, and because they all have probability $\pi_{\theta}(J)$, the Second Borel–Cantelli lemma then tells us that, with $P_{\theta}$–probability 1, the event $\{I_{i}=J\}$ happens infinitely often as $i\to\infty$. So, any interval that has nonzero mass under the design will keep coming back, over and over again, along the infinite sequence of trials.
As in the infinite sequences above, each time our particular interval $J$ reappears, it comes with the same two layers of probability given to it by the model: before seeing $I_{i}$, the event $\{\theta\in I_{i}\}$ has unconditional design-level probability $1-\alpha$, while after seeing $I_{i}=J$, the corresponding indicator $Z_{i}$ takes its fixed value in $\{0,1\}$. The key point is that these are probabilities at different conditioning levels of the same model; this does not claim that $P_{\theta}(\theta\in J)=1-\alpha$ for a fixed realized $J$. Rather, $1-\alpha$ is the pre-conditioning probability attached to each trial before revealing which interval value occurred.
Of course, in the usual continuous CI story (for example, t intervals for a normal mean), the distribution of $I(X)$ is itself continuous, and so the probability of seeing exactly the same endpoints twice is zero (under the model): every exact numerical interval $J$ has $\pi_{\theta}(J)=0$. In those models the Borel–Cantelli picture above is best read as a discrete intuition rather than a literal claim about recurrence of specific endpoints. The substantive long–run content of the design is still captured by the Bernoulli coverage indicators $Z_{i}$ and the law of large numbers, however–just not by the reappearance of any particular interval values themselves.
3.4 Stating the Obvious
The formalization above helps us see two important things. First, the statement that a given interval either covers the parameter or not is always true, regardless of whether it is the one we are about to construct, the one we just constructed, or the one we will construct a million samples from now. By definition, for the $i$th use of the confidence procedure, the model simultaneously implies both
@@ 3.4 Stating the Obvious
These identities hold both ex ante and ex post, and the only difference is whether we treat $X_{i}$ as random or condition on its realized value. In probabilistic terms, they are just two layers of the same law of iterated expectations:
@@
Reporting $P_{\theta}(Z_{\theta}(X_{i})=1)=1-\alpha$ amounts to working with a coarser information field (the design alone), whereas reporting $P_{\theta}(Z_{\theta}(X_{i})=1\mid X_{i})$ uses the finer $\sigma$-algebra generated by $X_{i}$. Regarding our interpretation of a single CI, choosing to report the degenerate conditional probability rather than the unconditional $P_{\theta}(Z_{\theta}(X_{i})=1)=1-\alpha$ is therefore a choice of conditioning level (i.e., of $\sigma$-algebra), not something dictated by the model itself.
@@
even though in each realized world $\omega$ the interval $I_{i}(\omega)$ has fixed endpoints. From the Borel-Cantelli aside above, we also know that in the finite-population setting, this probability will also attach to the same numerical interval an infinite number of times (which means if we make any ex ante, design-level claims about coverage, we will eventually make the claim about an actual interval we have already constructed). In both cases we can see that, under the model, coverage probability really is attaching to the intervals themselves–just not the intervals conditioned on themselves–and that the “either-or” reading ex post does not arise from some deep mathematical asymmetry between pre- and post-data uses of the model, but rather from the choice of one conditioning level—the finest one—over the others the model also happily supports.
@@ 4 Discussion
4.1 Can we have it both ways?
@@ 4.1.1 The degenerate as an interpretation
4.1.2 The degenerate as the only interpretation
@@
4.2 What is confidence, then?
@@
4.3 A suggestion for ex post probability statements
@@
More straightforwardly, and perhaps a bit provocatively, I would like to suggest that as a matter of mathematical bookkeeping, we might reevaluate the standard frequentist notion of probability as living only in the process of random sampling. The “pre–data” and “post–data” distinction already has an epistemic flavor, being naturally defined by someone’s having observed sampling to have taken place (otherwise, how could we tell between the two?), and it suggests that probability under the given model has somehow vanished once sampling has occurred. That implication is not required by frequentism per se, at least as realized in the Kolmogorov axioms: a Bernoulli–distributed outcome $Y_{i}$ determined by data $X_{i}$ carries an unconditional probability of success $p$ under the model, regardless of whether it has yet been realized. Instead, it seems more faithful to the frequentist program to ground our probability statements purely in the $\sigma$–algebras given to us by the model, choosing among them as needed so as to make our resulting inferences as accurate and as coherent as possible, and to resist the temptation to identify probability entirely with whatever “chanciness” may be inherent in physical, real–world processes, including a statistician drawing random samples from an actual population. I would argue that the latter picture is more properly viewed as a kind of frequentism–propensity–theory hybrid [8, 18], and that we might not gain much in return by blurring those lines (if anything keeping them separate might help us see more clearly the pragmatic and philosophical benefits of both). On the other hand, hewing to a model-based view puts us directly in touch with the mathematical support we need to make precisely the sort of objective, long–run probability statements that frequentism was originally conceived to secure.

### 17 Mar 2026 v4

@@ 1.2 Paper overview and contributions
1.3 Disclaimers
1.3.1 Generative AI usage
To save time in developing the manuscript, OpenAI’s GPT 5.2 Thinking (extended thinking) was used to typeset equations in LaTeX and to conduct a preliminary literature review. To improve readability and ensure correctness, the same model was also used to check the final version of the manuscript for typographical and mathematical errors.
1.3.2 Data availability statement
Because no data were used to produce the results in this manuscript, no data has been made available.
1.3.3 CDC author disclaimer
The findings and conclusions in this report are those of the author and do not necessarily represent the official position of the Centers for Disease Control and Prevention.
@@ 3.1 Confidence intervals and coverage indicators
3.2 Infinite sequences of experiments
@@
3.3 Connection to the Law of Iterated Expectation
The formalization above helps clarify two related points. First, for the $i$th use of the confidence procedure, the model simultaneously implies
@@ 3.3 Connection to the Law of Iterated Expectation
Thus the same model supports both the unconditional design-level statement and the degenerate conditional statement given the realized sample. In probabilistic terms, they are related by the law of iterated expectation:
@@
Reporting $P_{\theta}(Z_{\theta}(X_{i})=1)=1-\alpha$ amounts to working at a coarser information level, whereas reporting $P_{\theta}(Z_{\theta}(X_{i})=1\mid X_{i})$ uses the finer $\sigma$-algebra generated by $X_{i}$. In this sense, the familiar either-or reading corresponds to privileging the finest conditioning level, not to any failure of the model to support the unconditional design-based statement.
@@
even though in each realized world $\omega$ the interval $I_{i}(\omega)$ has fixed endpoints. In the finite-support setting, this observation becomes especially compelling: because the $Z_{i}$ are independent and each interval realization with positive mass under the design satisfies $P_{\theta}(X_{i}=x^{*})>0$ for every $i$, the Second Borel–Cantelli lemma guarantees that any such realization $x^{*}$ will recur infinitely often almost surely. That is, if we have already constructed an interval $I_{j}=I(x^{*})$ with particular numeric endpoints at some earlier index $j$, there will almost surely be a later index $i>j$ at which $X_{i}=x^{*}$ and $I_{i}=I(x^{*})$ exactly. At that later index, the model assigns the same design-level coverage probability as it always does:
$P_{\theta}\bigl(Z_{\theta}(X_{i})=1\bigr)=1-\alpha.$
But notice what this means: the observer already knows the numeric values of the endpoints that $I_{i}$ will take, having seen them when $I_{j}$ was constructed, and yet the model’s probability assignment is unchanged. Merely knowing the interval’s endpoints—the very information that the ”either-or” reading treats as collapsing coverage to $\{0,1\}$—does not, and cannot, alter the design-level probability that the model attaches to the coverage event. What this suggests is that the familiar post-data collapse to $\{0,1\}$ is not forced on us by anything intrinsic to having observed the data; rather, it reflects a choice to condition on the finest available $\sigma$-algebra, when the model equally supports the coarser one under which coverage remains $1-\alpha$.
@@ 4 Discussion
4.1 Competing ex post interpretations
@@ 4.1.1 The degenerate conditional as one interpretation
4.1.2 The degenerate conditional as the only interpretation
@@
4.2 Toward an interpretation of confidence
@@
4.3 A proposal for ex post probability statements
@@
More straightforwardly, and perhaps a bit provocatively, I would like to suggest that as a matter of mathematical bookkeeping, we might reevaluate the standard frequentist notion of probability as living only in the process of random sampling. The “pre–data” and “post–data” distinction already has an epistemic flavor, being naturally defined by someone’s having observed sampling to have taken place (otherwise, how could we tell between the two?), and it suggests that probability under the given model has somehow vanished once sampling has occurred. That implication is not required by frequentism per se, at least as realized in the Kolmogorov axioms: a Bernoulli–distributed outcome $Y_{i}$ determined by data $X_{i}$ carries an unconditional probability of success $p$ under the model, regardless of whether it has yet been realized. Instead, it seems more faithful to the frequentist program to ground our probability statements purely in the $\sigma$–algebras given to us by the model, choosing among them as needed so as to make our resulting inferences as accurate and as coherent as possible, and to resist the temptation to identify probability entirely with whatever “chanciness” may be inherent in physical, real–world processes, including a statistician drawing random samples from an actual population. I would argue that the latter picture is more properly viewed as a kind of frequentism–propensity–theory hybrid [8, 18], and that we might not gain much in return by blurring those lines (if anything keeping them separate might help us see more clearly the pragmatic and philosophical benefits of both). On the other hand, hewing to a model-based view puts us directly in touch with the mathematical support we need to make precisely the sort of objective, long–run probability statements that frequentism was originally conceived to secure..
5 Acknowledgments
Thanks to Chad Heilig for helping me refine the mathematical ideas for this paper; to Mariette Marano-Lee for helping me refine the thought experiments; and to Cheri Gatland-Lightner for carefully reviewing the final draft.